# The study of coherent Rayleigh-Brillouin scattering in multiple flow regimes using unified gas-kinetic scheme

### v1

@@ 1 Introduction
Due to molecular motion, light undergoes scattering as it propagates through a gas. In most cases, the resulting scattered spectrum exhibits multiple peaks, each associated with a distinct scattering mechanismyoung1982rayleigh, grinstead2000coherent, wu2022rarefied. This feature is particularly evident in Rayleigh–Brillouin Scattering (RBS), whose spectrum primarily comprises two components: Rayleigh and Brillouin scattering. Rayleigh scattering is a linear elastic process arising from Doppler shifts induced by the thermal motion of individual gas molecules and exhibits a much larger intrinsic cross-section than inelastic processes such as Raman scattering. In contrast, Brillouin scattering is a collective process involving acoustic phonons within the gas. The coherent molecular motion associated with these phonons behaves as a set of “macroscopic scatterers”, concentrating the energy and photon distribution of the scattered light and thereby enhancing the effective scattering cross-sectionmiles2001laser, 10.1063/5.0151530. Owing to its relatively large scattering cross-section, RBS enables the retrieval of macroscopic fluid parameters through rapid spectral measurement and fitting, making it a powerful non-destructive optical diagnostic technique for gas characterization10.1063/5.0268782, 10.1063/5.0225755, 10.1063/5.0226162, 10.1063/5.0129353. This method has found broad applications in meteorological observation and atmospheric remote sensingzhai2020rayleigh, pi2025rapid, hypersonic wind tunnel testingpanda2020spectrally, plasma diagnosticsshneider2007molecular, barker2001optical, and investigations of combustion kineticstrueba2025filtered.
@@
Building on its application in fluid parameter measurements, researchers have developed and investigated various forms of Rayleigh–Brillouin Scattering (RBS), notably spontaneous RBS (SRBS) and coherent RBS (CRBS). In SRBS, incident light with a given wave vector is scattered due to spontaneous fluctuations in gas densitysrbswangphd, 10.1063/5.0276580, 10.1063/5.0228567, ma2021molecular. Conversely, CRBS operates through the generation of periodic refractive-index gratings within a medium. These gratings are produced by the interference of two pump laser beams, which form a spatially periodic light-intensity pattern inside the medium. The physical mechanism of CRBS in gases is illustrated in Fig.1. By linearly varying the frequency difference between the two pump beams, interference patterns with different velocities can be created. Under the influence of the ponderomotive force, gas particles accumulate in the high-intensity regions of the interference pattern, forming an optical lattice. When a third beam—referred to as the probe beam—is incident on this lattice at an angle satisfying the Bragg condition, it undergoes coherent scattering, generating a signal beam. Measuring the signal beam’s intensity as a function of the lattice velocity yields the RBS spectrumPhysRevLett.89.183001. The CRBS spectrum typically exhibits a triplet structure, comprising a Rayleigh peak centered at zero frequency difference and broadened by Doppler effects due to molecular thermal motion, along with two Brillouin side peaks symmetrically positioned about the Rayleigh peak. These side peaks, centered near the medium’s sound speed, result from collective acoustic fluctuations. Unlike SRBS, where scattering occurs isotropically, CRBS confines the scattered light to a specific direction through phase matching, thereby achieving a much higher signal-to-noise ratio. In the kinetic regime, the CRBS line shape serves as a powerful diagnostic tool for probing gas properties and exploring microscopic kinetic processes.
@@
Among the existing numerical studies of Coherent Rayleigh–Brillouin Scattering (CRBS), several kinetic line-shape models for Maxwellian molecules have been developed, such as the Tenti S6 modeltenti1974kinetic and Pan’s S6/S7 modelsPhysRevLett.89.183001. These models are derived from simplified linearized Boltzmann equations—specifically, the linearized Bhatnagar–Gross–Krook (BGK) equation for atomic gases and the Wang–Chang–Uhlenbeck equation for molecular gases—and have shown good agreement with experimental observations. Under weaker modeling assumptions, the Direct Simulation Monte Carlo (DSMC) method has also been applied to CRBS systems both withsuzuki2024effects and withoutcornella2012experimental chirped lasers to simulate the rarefied flow regime. However, DSMC is computationally expensive due to inherent statistical noise, especially for unsteady problems. Moreover, because DSMC decouples molecular free transport and collisions, its spatial and temporal resolutions are constrained by the molecular mean free path and mean collision time, respectively, leading to substantially higher computational cost in the near-continuum regime. In this work, the deterministic Unified Gas-Kinetic Scheme (UGKS)xu2010unified is extended for CRBS simulations. By directly modeling the underlying physical laws within a discretized space, UGKS transcends the microscopic scale limitations typical of near-continuum flows. Specifically, its integral solution of the kinetic model is used to construct fluxes that couple free transport and collision effects of the gas distribution function, recovering the Gas-Kinetic Scheme (GKS)xu2001gas in the macroscopic limit. The UGKS has been successfully applied to a range of physical systems, including subsonic flows10.1063/5.0216358, hypersonic flowslong_implicit_2024, thermal and chemical non-equilibrium flowswei2024adaptive, wei2025unified, radiation transportquan2025radiative, plasma transferliu2017unified, and multiphase systemsLIU2019264. Here, UGKS is employed to simulate CRBS using the Strang splitting method to handle the perturbation source term. Compared with DSMC, UGKS allows for larger time steps and cell sizes in near-continuum regimes and achieves higher computational efficiency by eliminating particle noise. On the other hand, simulation efficiency is enhanced through the Linear Boltzmann Equation (LBE) formulation and Fourier-domain treatment, solved using the fast spectral methodwu2015kinetic and the general synthetic iteration scheme (GSIS)su2020fast. Their results show excellent agreement with both experimental measurements and DSMC simulations, and the approach has been successfully extended to diatomic gaseswu2025coherent. While the current method offers significantly higher efficiency than conventional numerical solvers, the key advantage of UGKS lies in its capacity to model larger perturbation accelerations arising from source terms by employing a BGK-type kinetic model instead of a linearized one. Accordingly, this study explores the influence of strong perturbations across different Knudsen numbers (Kn), extending beyond the conventional small-perturbation assumption and addressing physical scenarios similar to those discussed in Refs.pan_coherent_2003, shneider2007molecular, barker2001optical.

### v2

@@ 1 Introduction
Due to molecular motion, light undergoes scattering as it propagates through a gas. In most cases, the resulting scattered spectrum exhibits multiple peaks, each associated with a distinct scattering mechanismyoung1982rayleigh, grinstead2000coherent, wu2022rarefied. This feature is particularly evident in Rayleigh–Brillouin Scattering (RBS), whose spectrum primarily comprises two components: Rayleigh and Brillouin scattering. Rayleigh scattering is a linear elastic process arising from Doppler shifts induced by the thermal motion of individual gas molecules and exhibits a much larger intrinsic cross-section than inelastic processes such as Raman scattering. In contrast, Brillouin scattering is a collective process involving acoustic phonons within the gas. The coherent molecular motion associated with these phonons behaves as a set of “macroscopic scatterers”, concentrating the energy and photon distribution of the scattered light and thereby enhancing the effective scattering cross-sectionmiles2001laser. Owing to its relatively large scattering cross-section, RBS enables the retrieval of macroscopic fluid parameters through rapid spectral measurement and fitting, making it a powerful non-destructive optical diagnostic technique for gas characterization. This method has found broad applications in meteorological observation and atmospheric remote sensingzhai2020rayleigh, pi2025rapid, hypersonic wind tunnel testingpanda2020spectrally, plasma diagnosticsshneider2007molecular, barker2001optical, and investigations of combustion kineticstrueba2025filtered.
@@
Building on its application in fluid parameter measurements, researchers have developed and investigated various forms of Rayleigh–Brillouin Scattering (RBS), notably spontaneous RBS (SRBS) and coherent RBS (CRBS). In SRBS, incident light with a given wave vector is scattered due to spontaneous fluctuations in gas densitysrbswangphd, park2025physics, zhang2009monte, ma2021molecular. Conversely, CRBS operates through the generation of periodic refractive-index gratings within a medium. These gratings are produced by the interference of two pump laser beams, which form a spatially periodic light-intensity pattern inside the medium. The physical mechanism of CRBS in gases is illustrated in Fig.1. By linearly varying the frequency difference between the two pump beams, interference patterns with different velocities can be created. Under the influence of the ponderomotive force, gas particles accumulate in the high-intensity regions of the interference pattern, forming an optical lattice. When a third beam—referred to as the probe beam—is incident on this lattice at an angle satisfying the Bragg condition, it undergoes coherent scattering, generating a signal beam. Measuring the signal beam’s intensity as a function of the lattice velocity yields the RBS spectrumPhysRevLett.89.183001. The CRBS spectrum typically exhibits a triplet structure, comprising a Rayleigh peak centered at zero frequency difference and broadened by Doppler effects due to molecular thermal motion, along with two Brillouin side peaks symmetrically positioned about the Rayleigh peak. These side peaks, centered near the medium’s sound speed, result from collective acoustic fluctuations. Unlike SRBS, where scattering occurs isotropically, CRBS confines the scattered light to a specific direction through phase matching, thereby achieving a much higher signal-to-noise ratio. In the kinetic regime, the CRBS line shape serves as a powerful diagnostic tool for probing gas properties and exploring microscopic kinetic processes.
@@
Among the existing numerical studies of Coherent Rayleigh–Brillouin Scattering (CRBS), several kinetic line-shape models for Maxwellian molecules have been developed, such as the Tenti S6 modeltenti1974kinetic and Pan’s S6/S7 modelsPhysRevLett.89.183001. These models are derived from simplified linearized Boltzmann equations—specifically, the linearized Bhatnagar–Gross–Krook (BGK) equation for atomic gases and the Wang–Chang–Uhlenbeck equation for molecular gases—and have shown good agreement with experimental observations. Under weaker modeling assumptions, the Direct Simulation Monte Carlo (DSMC) method has also been applied to CRBS systems both withsuzuki2024effects and withoutcornella2012experimental chirped lasers to simulate the rarefied flow regime. However, DSMC is computationally expensive due to inherent statistical noise, especially for unsteady problems. Moreover, because DSMC decouples molecular free transport and collisions, its spatial and temporal resolutions are constrained by the molecular mean free path and mean collision time, respectively, leading to substantially higher computational cost in the near-continuum regime. In this work, the deterministic Unified Gas-Kinetic Scheme (UGKS)xu2010unified is extended for CRBS simulations. By directly modeling the underlying physical laws within a discretized space, UGKS transcends the microscopic scale limitations typical of near-continuum flows. Specifically, its integral solution of the kinetic model is used to construct fluxes that couple free transport and collision effects of the gas distribution function, recovering the Gas-Kinetic Scheme (GKS)xu2001gas in the macroscopic limit. The UGKS has been successfully applied to a range of physical systems, including hypersonic flowslong_implicit_2024, jiang2019, micro flowsliuchang2020, thermal and chemical non-equilibrium flowswei2024adaptive, wei2025unified, radiation transportquan2025radiative, plasma transferliu2017unified, and multiphase systemsLIU2019264. Here, UGKS is employed to simulate CRBS using the Strang splitting method to handle the perturbation source term. Compared with DSMC, UGKS allows for larger time steps and cell sizes in near-continuum regimes and achieves higher computational efficiency by eliminating particle noise. On the other hand, simulation efficiency is enhanced through the Linear Boltzmann Equation (LBE) formulation and Fourier-domain treatment, solved using the fast spectral methodwu2015kinetic and the general synthetic iteration scheme (GSIS)su2020fast. Their results show excellent agreement with both experimental measurements and DSMC simulations, and the approach has been successfully extended to diatomic gaseswu2025coherent. While the current method offers significantly higher efficiency than conventional numerical solvers, the key advantage of UGKS lies in its capacity to model larger perturbation accelerations arising from source terms by employing a BGK-type kinetic model instead of a linearized one. Accordingly, this study explores the influence of strong perturbations across different Knudsen numbers (Kn), extending beyond the conventional small-perturbation assumption and addressing physical scenarios similar to those discussed in Refs.pan_coherent_2003, shneider2007molecular, barker2001optical.
@@ 5 Conclusions
Data availability
The data that supports the findings of this study are available within the article and its supplementary material.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.